# The Epistemic Support-Point Filter: Jaynesian Maximum Entropy Meets Popperian Falsification A Possibilistic Minimax-Entropy Optimality Proof

### 10 Mar 2026 v1

@@ 6.1 Setup
We validate the theorem’s regime structure numerically using the ESPF applied to a range and range-rate orbital tracking scenario. The filter operates in state dimension $n=7$ with Smolyak Level 3 quadrature, yielding $M=106$ support points and $N_{\mathrm{target}}\approx 81$8383 in nominal tracking, 32326464 at stress events. Process noise is $a_{R}=a_{T}=a_{N}=5\times 10^{-12}$ km/s2 in RTN coordinates. Measurement noise is $\sigma_{R}=1$ m range, $\dot{\sigma}_{R}=10^{-5}$ km/s range-rate. Ground stations: Arecibo, Kwajalein, Diego Garcia.
@@
The Epistemic Width Monitor (EWM) is computed at every step and comprises:
  • $W_{\mathrm{ep}}$: epistemic width (0 = collapsed/probabilistic, 1 = total ignorance),
  • $\log\det(\mathrm{MVEE}_{\lambda})$: regime indicator (negative = Popperian),
  • $\alpha_{c}$: Hölder exponent (regime-sensitive geometry),
  • $I_{c}$: possibilistic information content (Shannon surrogate proxy; see Remark 6.1).
@@
The EWM implementation computes $I_{c}$ via a Shannon entropy surrogate rather than the true $\alpha$-cut integral $H_{\pi}$. This proxy underestimates $I_{c}$ in the falsification regime where it matters most. The EWM machinery for the true numerical $\alpha$-cut integral exists in espf_lemma3_diagnostic and will replace the surrogate in a future release. The qualitative diagnostic value of $I_{c}$ as an early-warning indicator (see Section 6.4) is unaffected by this approximation.
@@ 6.2 Smolyak Level 3: Nominal Run
The nominal 2-day run demonstrates the filter operating entirely in the diffusion regime: $\log\det(\mathrm{MVEE})$ remains positive throughout ($\approx 100$120120 nats), and the regime flag never flips to Popperian. This is the correct behavior—a well-conditioned nominal LEO track with no model mismatch does not generate a falsification event.
@@
  • Prune count settles to 12121616 points per step after an initial 84-point contraction at the first measurement. This tight, stationary band confirms the filter has found its operating regime.
  • $\sigma_{k}$ stabilizes to a stationary band of $\approx 0.45$$0.85$ after initialization, with no systematic drift indicating filter divergence.
  • Necessity hugs zero with episodic flares to 0.020.020.060.06, and surprisal maintains a stationary band of $\approx 0.07$$0.13$. Flat surprisal is a model-fidelity indicator: the filter is not progressively surprised by its own predictions.
  • $\alpha_{c}$ rides a stable band near 2.12.12.32.3 with periodic measurement-epoch spikes to $\approx 3.0$, reflecting momentary near-falsification-pressure that resolves without a full regime flip.
The Smolyak Level 3 nominal run confirms what the theorem predicts: under clean conditions, the filter operates in the Jaynesian regime, and all EWM indicators are stationary.
@@ 6.3 Smolyak Level 3: Stress Run
EWM regime indicator: $\log\det(\mathrm{MVEE})$ remains positive throughout the 2-day run ($\approx 80$120120 nats). The MVEE regime flag never flips to Popperian. This is a significant finding discussed below.
Prune count (the key diagnostic): Before day 0.75, the prune count matches the nominal pattern (12121616 points per step). At day $\approx 0.75$, coinciding with the accumulated effect of the range bias and approaching the maneuver epoch, the prune count escalates sharply to 40408585 points per step and remains elevated for the remainder of the run. At $M=106$, this represents 38388080% of the support cloud being pruned at each update step. A prominent horizontal stripe at 42 pruned points reflects a grid-boundary constraint in the Smolyak point configuration; this is a finite-grid artifact and does not affect the diagnostic interpretation.
Necessity and surprisal (the early-warning indicators): Before day 0.75, necessity is near zero and surprisal is in the nominal band (0.050.050.120.12). Beginning at day $0.75$, necessity escalates dramatically, with values routinely 0.40.41.01.0 from day $1.0$ onward. Necessity approaching 1.0 means individual support points are becoming necessary: the measurement is inconsistent with the support if they are removed. The surprisal axis, scaled to $[0,1000]$ compared to the nominal $[0,0.2]$, shows peaks of $\approx 640$ (day 1.0), $\approx 890$ (day 1.8), and episodic spikes >100>100 throughout. A surprisal of 900 means the filter is encountering a measurement that is nearly outside its predictive support.
@@ 6.4 EWM Diagnostic Findings
Comparison with Level 2. At Smolyak Level 2 ($M=N_{\min}$), the prune count band is 12121616 in both nominal and stress cases, and necessity/surprisal diverge at the stress onset but with smaller magnitude. The MVEE regime flag never flips at Level 2 either. This confirms that the regime structure is a property of the underlying theory—not a grid-resolution artifact—and that Level 3 provides diagnostic resolution that Level 2 cannot: specifically, the escalation in prune count from 15 to 80 at Level 3 is invisible at Level 2 where the grid is already at minimum cardinality.
@@ 6.5 Tabular Validation (52 Steps)
Tables 1 and 2 present the full 52-step Claim A/B diagnostic from the stress run, validating Lemma 3.6 directly. Key findings from this record:
@@
Table 2: Lemma 3 diagnostic, steps 19–52: deep contraction, second collapse, re-stabilization. $N_{t}=N_{\mathrm{target}}$; $\star$ marks the second stress event.
@@
[…]
@@ 7.1 What the Theorem Establishes and What Remains Open
The one outstanding proof obligation is Conjecture 7.1.
Conjecture 7.1 (Isotropy for orbital mechanics).
For range and range-rate measurements of trajectories propagated under the two-body problem or its low-order perturbations, the innovation-state isotropy condition (17) holds to first order in the linearization residual over observation intervals short relative to the orbital period.
For the linear-measurement case and for locally linearizable nonlinear models satisfying condition (17), the theorem is complete.
@@ 7.3 Comparison to Existing Optimality Results
Filter Optimality criterion Assumptions
Kalman Min. posterior covariance (MMSE) Linear, Gaussian
Particle Asymptotically exact posterior Non-Gaussian, stochastic
ESPF Min. possibilistic entropy $H_{\pi}$ Bounded, possibilistic
@@ 7.4 Implications for Filter Design
Possibility assignment is not optional. A filter that selects survivors by minimum $q^{(i)}_{k}$ but reassigns uniform possibility forfeits the gradient entropy improvement. The compatibility-based assignment $\rho^{\star}(i)=\mathrm{Comp}^{(i)}_{k}$ is constitutive of the ESPF’s optimality, not a refinement.
@@ 8 Conclusion
The ESPF was designed around a conviction: that a rational recursive estimator should be quick to embrace ignorance and slow to assert certainty. This paper has shown that this conviction is a theorem.

### 12 Mar 2026 v2

@@ 6.1 Setup
We validate the theorem’s regime structure numerically using the ESPF applied to a range and range-rate orbital tracking scenario. The filter operates in state dimension $n=7$ with Smolyak Level 3 quadrature, yielding $M=106$ support points and $N_{\mathrm{target}}\approx 81$–83 in nominal tracking, 32–64 at stress events. Process noise is $a_{R}=a_{T}=a_{N}=5\times 10^{-12}$ km/s2 in RTN coordinates. Measurement noise is $\sigma_{R}=1$ m range, $\sigma_{\dot{R}}=10^{-5}$ km/s range-rate. Ground stations: Arecibo, Kwajalein, Diego Garcia.
@@
The Epistemic Width Monitor (EWM) is computed at every step and comprises: $W_{\mathrm{ep}}$ (epistemic width), $\log\det(\mathrm{MVEE}_{\lambda})$ (regime indicator), $\alpha_{c}$ (Hölder exponent), and $I_{c}$ (possibilistic information content; see Remark 6.1).
@@
The current EWM implementation computes $I_{c}$ via a Shannon entropy surrogate rather than the true possibilistic $\alpha$-cut integral $H_{\pi}$. This proxy underestimates $I_{c}$ in the falsification regime where it matters most; a full numerical implementation of the $\alpha$-cut integral is planned for a future release. In the stress run, the surrogate underestimates $I_{c}$ by a factor of 2 or more in high-surprisal epochs (days 1.0–1.8); the qualitative escalation pattern is preserved but the absolute magnitudes reported in Section 6.4 should be treated as lower bounds on true possibilistic stress.
@@ 6.2 Smolyak Level 3: Nominal Run
The nominal 2-day run demonstrates the filter operating entirely in the diffusion regime: $\log\det(\mathrm{MVEE})$ remains positive throughout ($\approx 100$–120 nats), and the regime flag never flips to Popperian. This is the correct behavior—a well-conditioned nominal LEO track with no model mismatch does not generate a falsification event.
@@
  • Prune count settles to 12–16 points per step after an initial 84-point contraction at the first measurement.
  • $\sigma_{k}$ stabilizes to a stationary band of $\approx 0.45$–0.85 after initialization, with no systematic drift indicating filter divergence.
  • Necessity hugs zero with episodic flares to 0.02–0.06, and surprisal maintains a stationary band of $\approx 0.07$–0.13.
  • $\alpha_{c}$ rides a stable band near 2.1–2.3 with periodic measurement-epoch spikes to $\approx 3.0$.
@@ 6.3 Smolyak Level 3: Stress Run
EWM regime indicator: $\log\det(\mathrm{MVEE})$ remains positive throughout the 2-day run ($\approx 80$–120 nats). The MVEE regime flag never flips to Popperian. This is a significant finding discussed below.
Prune count (the key diagnostic): Before day 0.75, the prune count matches the nominal pattern (12–16 points per step). At day $\approx 0.75$, coinciding with the accumulated effect of the range bias and approaching the maneuver epoch, the prune count escalates sharply to 40–85 points per step and remains elevated for the remainder of the run. At $M=106$, this represents 38–80% of the support cloud being pruned at each update step.
Necessity and surprisal (the early-warning indicators): Before day 0.75, necessity is near zero and surprisal is in the nominal band (0.05–0.12). Beginning at day 0.75, necessity escalates dramatically, with values routinely 0.4–1.0 from day 1.0 onward. The surprisal axis, scaled to $[0,1000]$ compared to the nominal $[0,0.2]$, shows peaks of $\approx 640$ (day 1.0), $\approx 890$ (day 1.8), and episodic spikes >100>100 throughout.
@@ 6.4 EWM Diagnostic Findings
Comparison with Level 2. At Smolyak Level 2 ($M=N_{\min}$), the prune count band is 12–16 in both nominal and stress cases, and necessity/surprisal diverge at the stress onset but with smaller magnitude. The MVEE regime flag never flips at Level 2 either. This confirms that the regime structure is a property of the underlying theory—not a grid-resolution artifact—and that Level 3 provides diagnostic resolution that Level 2 cannot.
@@ 6.5 Tabular Validation (52 Steps)
Tables 1 and 2 present the full 52-step Claim A/B diagnostic from the stress run, validating Lemma 3.7 directly.
Key findings:
@@
Table 2: Lemma 3 diagnostic, steps 19–52: deep contraction, second collapse, re-stabilization. $N_{t}=N_{\mathrm{target}}$; marks the second stress event.
@@
Step 35: second stress event. $N_{t}=52$; min-$q$ beats random by factor $\sim 11$.
@@ 7.1 What the Theorem Establishes and What Remains Open
For the linear-measurement case and for locally linearizable nonlinear models satisfying condition (10), the theorem is complete. Proposition 3.10 establishes the isotropy condition for the orbital mechanics setting to first order; a full proof of the nonlinear isotropy bound across all observation geometries remains open.
@@ 7.3 Comparison to Existing Optimality Results
Filter Optimality criterion Assumptions Filter is for
Kalman Min. posterior covariance (MMSE) Linear, Gaussian Estimating truth
Particle Asymptotically exact posterior Non-Gaussian, stochastic Estimating truth
ESPF Min. possibilistic entropy $H_{\pi}$ Bounded, possibilistic Eliminating the impossible
The fourth column is the deepest distinction. Bayesian filters — Kalman and particle alike — are driven by a model of how truth generates observations. Their optimality guarantees are statements about convergence to truth under that model. When the model is wrong, they converge to the wrong truth. The ESPF carries no model of truth. Its optimality guarantee is a statement about ignorance: it retains the minimum set of hypotheses consistent with the evidence, assigns them possibilities in proportion to their evidence compatibility, and makes no further claim. It cannot be wrong about truth because it never asserts truth. It can only be more or less efficient at eliminating the impossible — and the EWM measures exactly that efficiency in real time.
@@
The Hölder mean structure (Remark 5.3) makes this hierarchy precise and connects it directly to the numerical findings. The three optimality criteria occupy positions $p=+\infty$, $p=0$, and $p=0$ (Gaussian) on the Hölder curve. The ordering $M_{0}\leq M_{+\infty}$ explains a finding that might otherwise seem puzzling: Claim B ($H^{\omega}_{\pi}$, order 0) is more robustly satisfied than Claim A ($\log\det\mathrm{MVEE}$, order $\infty$) throughout the tabular validation, including at steps where Claim A shows swap failures. This is not a numerical artifact. It is a direct consequence of the Hölder ordering: the geometric mean of $\alpha$-cut volumes is intrinsically less sensitive to perturbation at the boundary of the support than the supremum. The MVEE regime flag, operating at $p=+\infty$, is the coarsest instrument in the EWM suite for exactly this reason. The ESPF’s optimality target $H_{\pi}$ is the more stable, more informative criterion — and the numerical results confirm it.
@@ 7.4 Implications for Filter Design
Possibility assignment is not optional. A filter that selects survivors by minimum $q^{(i)}_{k}$ but reassigns uniform possibility forfeits the gradient entropy improvement. The compatibility-based assignment $\rho^{*}(i)=\mathrm{Comp}^{(i)}_{k}$ is constitutive of the ESPF’s optimality, not a refinement.
@@ 8 Conclusion
This paper has proved that the Epistemic Support-Point Filter (ESPF) is the unique optimal recursive estimator within the class of epistemically admissible evidence-only filters. Where Bayesian filters are driven toward an assumed truth, the ESPF eliminates the impossible: it retains only what the evidence has not ruled out, assigns possibilities in proportion to evidence compatibility, and makes no further claim. The prior has no vote on what is impossible. Only the evidence does.